The shells fired from an artillery piece have a muzzle speed of 195 m/s, and the target is at a horizontal distance of 2.00 km. Two possible angles can be used to hit the target. Find the smallest angle relative to the horizontal at which the gun should be aimed? (Hint: sin2θ = 2sinθcosθ)
Answer #1

Given R=2km=2000m

speed=195m/s

smallest angle Θ = ½*arcsin[R*g/V²] = 15.51° above horizontal

The larger angle Θ' = (90 - Θ) = 74.48°